A 8.0 V potential difference is applied between the ends of a 0.80-mm-diameter, 90-cm-long nichrome wire. What is the current in the wire?
Answer #1

Area

A=pi*D2/4 =pi*(0.8*10-3)2/4 =5.0265*10-4m2

Resistance of the wire is

R=pL/A =(1.5*10-6)(0.9)/(5.0265*10-4)

R=2.68574 ohms

Current flowing in the wire is

I=V/R =8/2.68574

I=2.9787 A